The generic metal hydroxide M(OH)2 has Ksp = 5.85×10−18.

A. What is the solubility of M(OH)2 in pure water?

B. What is the solubility of M(OH)2 in a 0.202 M solution of M(NO3)2?
Answer #1

A. ksp = 4s^3   = 5.85*10^(-18)

s = solubility = 1.135*10^(-6) M

B. Ksp = s(2s)^2 = 5.85*10^(-18)

           0.202*(2s)^2 = 5.85*10^(-18)

s = solubility = 2.69*10^(-9) M